From 3 rations of C2H2:O2, determine which has the most complete combustion reaction using balanced chemical reactions. The three ratios are: 100:0, 5:1 and 1.7:1
1
Expert's answer
2015-11-09T08:35:38-0500
Equilibrium of combustion reaction:
2C2H2 + 5O2 → 4CO2 + 2H2O
We need to have more oxygen for the same quality of the acetylene for more complete combustion.
We have three ratios:
The first ratio is 100:0. Here we have no oxygen. The combustion is impossible.
The second and third ratios have acetylene and oxygen: 5:1 and 1.7:1. Divide the both ratios into share of acetylene:
5:1 = 5/5:1/5 = 1:0.2
1.7:1 = 1.7/1.7 : 1/1.7 = 1:0.6
Third ratio has more complete combustion than second ration, because third ratio has more oxygen for the same quality of the acetylene.

Answer: third ratio.